The position of a particle moving along the x-axis is given by 

x = 3.0t2 - 2.0t3 where x is in meters and t is in seconds. What is the position 

of the particle when it achieves its maximum speed in the positive x-direction?


I think the answer is 1 m but in the solutions manual there is 0.50 m.

Please tell me why 1 m is wrong. If we use a graphing calculator we will see that maximum point at 1 s and its value is 1 m.




1
Expert's answer
2021-10-31T18:09:34-0400

The velocity of the particle is given as follows:


"v(t) = x'(t) = 6t - 6t^2"

The maximum value of velocity is achieved when its derivative is equal to 0:


"v'(t) = 6-12t_{max}=0\\
t_{max} = 0.5s"

The graph of "v(t)" confirms that indeed, velocity has maximum at 0.5s.




Substituting "t_{max}" to "x(t)", get:


"x(t_{max}) = 3\cdot 0.5^2 - 2\cdot 0.5^3 = 0.5m"



Answer. 0.5m